Let f(x)=x^3-6x^2+3x+10f(x)=x3−6x2+3x+10, then choose the set of correct options regarding f(x).

  1. If x \in [0, 1] \cup (10, \infty)x∈[0,1]∪(10,∞), then f(x)is positive.
  2. If x \in (-\infty, -1] \cup (3, 6)x∈(−∞,−1]∪(3,6), then f(x) is negative.
  3. If x \in (-\infty, 1] \cup (2, 5)x∈(−∞,1]∪(2,5), then f(x) is negative.
  4. If x \in [-2, 2] \cup (5, \infty)x∈[−2,2]∪(5,∞), then f(x) is positive.
  5. If x \in (-1, 2) \cup (5, \infty)x∈(−1,2)∪(5,∞), then f(x)is positive.




1
Expert's answer
2021-10-04T04:57:55-0400
"f(x)=0=>x^3-6x^2+3x+10=0"

"x^2(x+1)-7x(x+1)+10(x+1)=0"

"(x+1)(x^2-7x+10)=0"

"(x+1)(x-2)(x-5)=0"

"x_1=-1, x_2=2, x_3=5"

If "x<-1," then "f(x)<0."

If "-1<x<2," then "f(x)>0."

If "2<x<5," then "f(x)<0."

If "x>5," then "f(x)>0."


Answer:

5. If "x\in(-1, 2)\cup(5,\infin)," then "f(x)" is positive.